64 to the 7/9 power divided by 4 to the 3/4 power. Having trouble showing answer of 8.98.
Answer by jim_thompson5910(35256)

64 = 4^3


So

(64^7/9)/(4^3/4) = (4^(3*7/9))/(4^3/4)

(64^7/9)/(4^3/4) = (4^(21/9))/(4^3/4)

(64^7/9)/(4^3/4) = 4^(21/9- 3/4)

(64^7/9)/(4^3/4) = 4^(19/12)

(64^7/9)/(4^3/4) = 8.979696 ... Note: Use a calculator on this step.